What volume of a 0.354 M perchloric acid solution is required to neutralize 20.6 mL of a 0.193 M sodium hydroxide solution?

__________ mL perchloric acid
1
Expert's answer
2019-03-18T02:19:41-0400

HClO4 + NaOH -> NaClO4 + H2O

c=n/V, then n= c*V

n(NaOH) = 0.0206 L* 0.193 mol/L = 0.00398 mol

According to equation mole ratio n( HClO4):n(NaOH)= 1:1, then n(HClO4)=n(NaOH) = 0.00398 mol

c=n/V, then V= n/c = 0.00398 mol/0.354 mol/L = 0.0112 L= 11.2 mL

Answer: 11.2 mL